OpenStudy (crashonce):

A man sold an article for $x at a loss of 20% on the cost price. What would have bbeen the loss or gain percent if he sold it for 1.5x?

OpenStudy (crashonce):

@mathmath333

OpenStudy (mathmath333):

if he sold it for s.p.=\(80\) at a loss of \(20\%\) then c.p=\(100\) if new s.p=\(1.5\times 80=120\) \(\large\tt \begin{align} \color{black}{ \% change =\dfrac{120-100}{100}\times 100 }\end{align}\)

OpenStudy (mathmath333):

s.p.>c.p. so he will be in profit

OpenStudy (perl):

Doing this algebraically Let y = cost price profit = sale price - cost price so for the first sentence we are told that you sell the item at a loss of 20% on the cost price. This gives us the equation : profit = .8 * y - y = -.2y then it asks, what happens if you increase the selling price by 1.5x . since x = .8y 1.5x = 1.5*.8 y = 1.2y and 1.2 y - y = .20y. so there is a gain of 20% on the cost price.